# Are Humans as Brittle as Large Language Models?

### v1

@@ Abstract
The output of large language models (LLM) is unstable, due to both non-determinism of the decoding process as well as to prompt brittleness. While the intrinsic non-determinism of LLM generation may mimic existing uncertainty in human annotations through distributional shifts in outputs, it is largely assumed, yet unexplored, that the prompt brittleness effect is unique to LLMs. This raises the question: do human annotators show similar sensitivity to instruction changes? If so, should prompt brittleness in LLMs be considered problematic? One may alternatively hypothesize that prompt brittleness correctly reflects human annotation variances. To fill this research gap, we systematically compare the effects of prompt modifications on LLMs and identical instruction modifications for human annotators, focusing on the question of whether humans are similarly sensitive to prompt perturbations. To study this, we prompt both humans and LLMs for a set of text classification tasks conditioned on prompt variations. Our findings indicate that both humans and LLMs exhibit increased brittleness in response to specific types of prompt modifications, particularly those involving the substitution of alternative label sets or label formats. However, the distribution of human judgments is less affected by typographical errors and reversed label order than that of LLMs.
@@ 1 Introduction
Large language models (LLMs) have shown impressive capabilities in automatic data annotation tasks (Tan et al., 2024). However, practical applications are often hindered by variability in model outputs, leading to inconsistent predictions (Zhao et al., 2024; Stureborg et al., 2024). This variability encompasses both the inherent nondeterminism in probabilistic models (Song et al., 2025; Ouyang et al., 2025) as well as prompt brittleness, wherein minor changes in prompt phrasing leads to significant differences in outputs (Lu et al., 2022; Tang et al., 2024).
Of course, variability is also observed in human annotations. Annotators’ individual experiences, sociodemographic backgrounds, and moral values shape how they interpret and label content. Moreover, the uncertainty in human annotation behavior may increase with the inclusion of additional annotators, posing challenges for explanation and identification (Pavlick and Kwiatkowski, 2019). Traditionally, such disagreement among annotators has been treated as noise or bias (Milkowski et al., 2021), often resolved through techniques like majority voting to produce a single “gold” label (Sabou et al., 2014; Anand et al., 2024).
Recent research suggests that variability in human annotations should not be seen as a “problem” but as a reflection of diverse human interpretations (Plank, 2022). For example, in an emotion labeling task, the text ‘I ran into my partner when I got home at 2 a.m.’ may be labeled as joy or surprise, depending on the subjective interpretation.
As LLMs are trained to model human text, it is reasonable that their output distributions should somehow reflect the variability present in human annotations. However, while recent work has acknowledged the importance of human label variation, little attention has been paid to how this variation behaves across different instruction perturbations, particularly whether changes in annotator instructions affect the distribution of human responses analogously LLM’s prompt brittleness.
In this work, we draw a connection between human label variation and LLM predictions. Specifically, we investigate whether prompt brittleness is unique to LLMs, or if humans exhibit similar sensitivity to task instruction variations.
We propose a systematic method that formalizes prompt perturbation and investigates the distributional effects of such pertubations on both LLMs and human annotators. This method supports our exploration of the following research questions:
@@
The experimental repository and supplementary materials are available at https://www.uni-bamberg.de/en/nlproc/projects/inprompt/.
@@ 2.1.1 Prompt Brittleness
LLMs exhibit high sensitivity to variations in a prompt, even when changes are minimal (Brown et al., 2020; Jiang et al., 2020; Gao et al., 2021). This phenomenon is commonly referred to as prompt brittleness. Prior studies have examined a variety of prompt perturbations that can lead to substantial changes in model outputs.
One prominent issue is position bias, where LLMs exhibit preferences for labels appearing in specific positions within a prompt. Altering the order of answer options can substantially affect model behavior (Lu et al., 2022; Qiang et al., 2024; Wang et al., 2024a, c). Similarly, the position of the set of possible answers as a whole affects the output (Zheng et al., 2024).
Lexical perturbations such as typos or synonym substitutions also affect the output. While some findings suggest that LLMs are relatively robust to typographical errors (Movva et al., 2024; Wang et al., 2024b), other work highlights brittleness in response to paraphrased instructions, including variations in verbs, nouns, and conjunctions (Mizrahi et al., 2024). Ceron et al. (2024) explore the effect of instruction styles, comparing personal versus impersonal phrasing.
Various approaches exist to mitigate the effect of prompt variations, such as guiding the inference process (Yugeswardeenoo et al., 2024; Lampinen et al., 2022) or by requesting explanations (Ye and Durrett, 2022; Mishra et al., 2022). Similarly, eliciting verbalized confidence estimates can reduce calibration errors in model outputs (Tian et al., 2023).
Tjuatja et al. (2024) is the only study we are aware of which studies the relation to human variations. They demostrate that commercial LLMs exhibit sensitivity to prompt biases that elicit minimal change in human responses. However, their exploration of prompt modifications primarily relies on survey-style prompt variants and does not study human annotation variability. In contrast, our work systematically generalizes across a broader and more diverse set of prompt reformulations, aiming to reflect the correlation of variability between LLMs and human annotation behavior.
@@ 2.1.2 LLM Non-determinism
LLMs not only yield inconsistent outputs due to prompt brittleness, but also exhibit inherent non-determinism, even for stable prompts. Notably, even with a temperature of zero, variability persists when invoking LLM APIs such as ChatGPT (Ouyang et al., 2025). This stochastic behavior poses challenges for reproducibility in downstream applications.
Song et al. (2025) investigate the relationship between model size and output stability, finding that smaller LLMs may yield more consistent outputs when comparing the greedy decoding with different sampling strategies. To address non-determinism in practical (non-greedy) sampling scenarios, where LLMs produce a distribution of outputs for the same prompt, Hayes et al. (2025) propose probabilistic discoverable extraction. Their framework provides a principled way to evaluate LLM performance across a space of possible outputs, rather than relying on single-instance generation.
In this work, we enable the sampling decoding of LLMs by treating variability not as noise but as an intrinsic property of LLM output, to be analyzed on the same footing as human response variations.
@@ 2.2 Variability in Humans
Differences in human annotation are commonly treated as noise, often attributed to annotation errors (Zhang and de Marneffe, 2021). However, Plank (2022) argue that variability in human annotations can also arise from factors such as subjectivity, perspectivism, annotator disagreement, and the existence of multiple plausible views.
2.2.1 Subjectivity & Perspectivism
Subjective tasks, such as hate speech detection and emotion classification, are inherently influenced by the value systems, cognitive frameworks, and personal experiences of annotators (Sap et al., 2019; Aroyo and Welty, 2015; Milkowski et al., 2021). In the context of emotion analysis, subjectivity has been studied across various domains, including healthcare, hospitality, and social media, where accounting for individual differences has been shown to enhance model performance in capturing user emotions (Miłkowski et al., 2022; Suzuki et al., 2022; Kazienko et al., 2023).
Recent work has introduced this concept in humans to LLMs, e.g., with socio-demographic prompting (Schäfer et al., 2025; Dayanik et al., 2022). Our approach considers the distribution of predictions by prompting LLMs a large number of times without anchoring responses to predefined demographic categories.
2.2.2 Annotator Disagreement
While subjectivity and perspectivism highlight the inherent variability in human interpretation, annotator disagreement also points to varying annotations which might not be of any benefit. Disagreement is therefore also observed in tasks considered objective, demanding a unified consensus for both model training and evaluation (Uma et al., 2022).
For example, Plank et al. (2014) demonstrate that disagreements in part-of-speech tagging stem from linguistically debatable cases rather than annotator mistakes. Similarly, Webber and Joshi (2012) analyze the challenges of human disagreement in discourse properties. Methods that go beyond simple majority voting address such issues. Das et al. (2017) propose to distinguish genuine annotation errors from legitimate disagreements in rhetorical structure analysis. Mostafazadeh Davani et al. (2022) advocate for multi-annotator modeling.
Semantic annotation tasks are also prone to disagreement. Sommerauer et al. (2020) evaluate such cases using multiple quality metrics. de Marneffe et al. (2012) argue that fact judgments should be treated as distributions. Pavlick and Kwiatkowski (2019) show that disagreements persist with amplified annotator numbers and contextual information.
@@ 3. How do you rate … given one of the labels ‘a’; ‘b’; ‘c’; ‘d’?
Table 1: Prompt modification methods in the neutral category. Modifications are bold in the text for each example.
@@ 3 Prompt Perturbation Methods
Sensitive changes might involve e.g. the introduction of emotional wording to the prompt (aspect name Emo). For this aspect, appending the sentence "Respond quickly!" to the prompt corresponds to the variant fast. Appendix A.3 provides the full list of evaluated prompts in our experiments, developed by this method.
@@ 4. How do you rate … given one of the labels ‘a’, ‘b’, ‘c’, ‘d’? Respond fast!
Table 2: Prompt modification methods in the sensitive category. Modifications are bold in the text for each example.
@@ 4 Experimental Settings
For both humans and LLMs, we consider four tasks from three distinct English-language datasets: (1) offensiveness rating and (2) politeness rating from the POPQUORN dataset (Pei and Jurgens, 2023), (3) irony detection from the EPIC dataset (Frenda et al., 2023), and (4) emotion classification from the crowd-enVent dataset (Troiano et al., 2023). These tasks were selected as all are known to involve some level of annotator subjectivity.
For each dataset, we construct a base prompt, which clearly and succinctly requests an annotation for a provided instance according to the task’s label set. We then construct a set of prompt variations, as described in Section 3. A full list of all prompt variations thus obtained is detailed in Appendix A.3.
@@
We select five LLMs capable of local deployment for evaluation: LLaMA-3.1-8B and LLaMA-3.3-70B (Patterson et al., 2022), Mixtral-8x7B (MistralAI, 2023), Falcon3-7B (TII, 2024), and Mistral-7B (Jiang et al., 2023). [^1] https://huggingface.co/: tiiuae/Falcon3-7B-Instruct, mistralai/Mixtral-8x7B-Instruct-v0.1, meta-llama/Llama-3.3-70B-Instruct, meta-llama/Llama-3.1-8B-Instruct.
@@
Model predictions are elicited in a zero-shot fashion: For each instance, the model is presented with one prompt variation, followed by the instance text, and allowed to generate freely. The generated text is then parsed by attempting to extract one of the requested labels, with unparsable responses discarded. Across all tasks, for each (prompt variation, instance, LLM) triple, we elicit 100 responses in this manner. We discard instances for which we obtain less than 100 total valid predictions after 500 attempts. This leaves us with final retained instances across all tasks, for which the statistics are shown in Table 3.
Human Study.
Due to the significantly higher costs associated with human annotations, we design our human study to collect a subset of the judgments collected in our LLM study. We conduct our human annotations on two tasks: offensiveness rating and emotion classification. For each task, we sample approximately 10% of instances from the original POPQUORN and crowd-enVent datasets, selecting an equal number of instances for each gold-standard label. We use five prompts per task for the human study, selected so as to maximize the diversity of prompts tested. In addition to the base prompt, we select the variants which caused the most and least extreme changes in LLM predictions from both the sensitive and neutral categories, as measured by Jensen-Shannon divergence (see Section 5). Each prompt-instance pair is re-annotated by 10 independent annotators recruited via the Prolific platform. [^2] https://www.prolific.com/ We employ only annotators based in the United States, with U.S. listed as their country of residence, country of birth, and nationality. All annotators are at least 18 years of age and demonstrated a high level of English proficiency, with English specified as their primary, first, and fluent language. We restrict the annotator approval rating to greater than 99%. Each annotator is presented with 22 instances including 2 attention checks, each preceded by a prompt variation acting as annotator instructions. To mitigate potential biases arising from prior exposure to the same task, each annotator is assigned only one prompt variant per task.
@@
Figure 1: Distance scores between the distributions of responses with prompt variants and the base prompt. The evaluated tasks are offensiveness rating (a) and emotion classification (b). The x-axis refers to the prompts with their modification types, and the y-axis refers to five LLMs and human samples. Distance scores are measured by Jensen-Shannon divergence (Dagan et al., 1997). Evaluated prompts can be found in Table LABEL:tab:prompt-off (a) and Table LABEL:tab:prompt-emo (b).

### v2

@@ Abstract
The output of large language models (LLMs) is unstable, due both to non-determinism of the decoding process as well as to prompt brittleness. While the intrinsic non-determinism of LLM generation may mimic existing uncertainty in human annotations through distributional shifts in outputs, it is largely assumed, yet unexplored, that the prompt brittleness effect is unique to LLMs. This raises the question: do human annotators show similar sensitivity to prompt changes? If so, should prompt brittleness in LLMs be considered problematic? One may alternatively hypothesize that prompt brittleness correctly reflects human annotation variances. To fill this research gap, we systematically compare the effects of prompt modifications on LLMs and identical instruction modifications for human annotators, focusing on the question of whether humans are similarly sensitive to prompt perturbations. To study this, we prompt both humans and LLMs for a set of text classification tasks conditioned on prompt variations. Our findings indicate that both humans and LLMs exhibit increased brittleness in response to specific types of prompt modifications, particularly those involving the substitution of alternative label sets or label formats. However, the distribution of human judgments is less affected by typographical errors and reversed label order than that of LLMs.
@@ 1 Introduction
Large language models (LLMs) have shown impressive capabilities in automatic data annotation tasks (tan-etal-2024-large). However, practical applications are often hindered by variability in model outputs, leading to inconsistent predictions (zhao-etal-2024-measuring; DBLP:journals/corr/abs-2405-01724). This variability encompasses both the inherent non-determinism in probabilistic models (song-etal-2025-good; 10.1145/3697010) as well as prompt brittleness, wherein minor changes in prompt phrasing lead to significant differences in outputs (lu-etal-2022-fantastically; tang-etal-2024-found).
Of course, variability is also observed in human annotations. Annotators’ individual experiences, sociodemographic backgrounds, and moral values shape how they interpret and label content. Moreover, the uncertainty in human annotation behavior may increase with the inclusion of additional annotators, posing challenges for explanation and identification (inherentdis). Traditionally, such disagreement among annotators has been treated as noise or bias (milkowski-etal-2021-personal), often resolved through techniques like majority voting to produce a single “gold” label (sabou-etal-2014-corpus; anand-etal-2024-dont).
Recent research suggests that variability in human annotations should not be seen as a “problem” but as a reflection of diverse human interpretations (plank-2022-problem). For example, in an emotion labeling task, the text ‘the dog ran towards me.’ may be labeled as joy or fear, depending on subjective preferences (troiano-etal-2023-dimensional).
As LLMs are trained to model human text, it is reasonable that their output distributions should somehow reflect the variability present in human annotations. However, while recent work has acknowledged the importance of human label variation, little attention has been paid to how this variation behaves across different prompt perturbations, particularly whether changes in annotator instructions affect the distribution of human responses analogously to LLM’s prompt brittleness.
In this work, we draw a connection between human label variation and LLM predictions. Specifically, we investigate whether prompt brittleness – a model’s sensitivity to small changes in task phrasing – is unique to LLMs, or if humans exhibit a comparable sensitivity to instruction variations. We propose a systematic method that formalizes prompt perturbation and investigates the distributional effects of such perturbations on both LLMs and human annotators. This method supports our exploration of the following research questions:
@@
The experimental repository and results are available at https://www.uni-bamberg.de/en/nlproc/projects/inprompt/.
@@ 2.1 Variability in LLMs
We focus on two factors for variability in LLMs: prompt brittleness and model non-determinism.
@@ 2.1.1 Prompt Brittleness
LLMs exhibit high sensitivity to variations in a prompt, even when changes are minimal (LMlearner; jiang-etal-2020-know; gao-etal-2021-making). This phenomenon is commonly referred to as prompt brittleness. Prior studies have examined a variety of prompt perturbations that can lead to substantial changes in model outputs.
One prominent issue is position bias, where LLMs exhibit preferences for labels appearing in specific positions within a prompt. Altering the order of answer options can substantially affect model behavior (lu-etal-2022-fantastically; qiang-etal-2024-prompt; wang-etal-2024-large-language-models-fair; wang-etal-2024-answer-c). Similarly, the position of the set of possible answers as a whole affects the output (zheng2024large).
Lexical perturbations such as typos or synonym substitutions also affect the output. While some findings suggest that LLMs are relatively robust to typographical errors (movva-etal-2024-annotation; wang2024look), other work highlights brittleness in response to paraphrased instructions, including variations in verbs, nouns, and conjunctions (mizrahi-etal-2024-state). ceron-etal-2024-beyond explore the effect of instruction styles, comparing personal versus impersonal phrasing.
Various approaches exist to mitigate the effect of prompt variations, such as guiding the inference process (yugeswardeenoo-etal-2024-question; lampinen-etal-2022-language) or by requesting explanations (Xprompt; mishra-etal-2022-cross). Similarly, eliciting verbalized confidence estimates can reduce calibration errors in model outputs (tian-etal-2023-just).
In this work, we will propose methods of prompt perturbations which build upon the theoretical foundations laid out in these works.
@@ 2.1.2 LLM Non-determinism
LLMs not only yield inconsistent outputs due to prompt brittleness, but also exhibit inherent non-determinism. Notably, even with a temperature of zero, variability persists when invoking LLM APIs such as ChatGPT (10.1145/3697010). This stochastic behavior poses challenges for reproducibility in downstream applications.
song-etal-2025-good investigate the relationship between model size and output stability, finding that smaller LLMs may yield more consistent outputs when comparing the greedy decoding with different sampling strategies. To address non-determinism in practical (non-greedy) sampling scenarios, where LLMs produce a distribution of outputs for the same prompt, hayes-etal-2025-measuring propose probabilistic discoverable extraction.
In this work, we employ non-deterministic sampling-based decoding strategies for LLMs, treating variability not as noise but as an intrinsic property of LLM output, to be analyzed on the same footing as human response variations.
@@ 2.2 Variability in Humans
In the context of human annotation, variations between annotators are commonly treated as noise, often attributed to annotation errors (zhang-de-marneffe-2021-identifying). However, plank-2022-problem argues that variability in human annotations can also arise from factors such as subjectivity, perspectivism, annotator disagreement, and the existence of multiple plausible views. Conversely, in survey responses, inter-participant variability is expected and desired in order to reflect the diversity of the surveyed population. However, such responses may also be sensitive to how questions are asked (survey-review), exhibiting similarities to the prompt-brittleness phenomenon observed for LLMs.
2.2.1 Subjectivity & Annotator Disagreement
Subjective tasks, such as hate speech detection and emotion classification, are inherently influenced by the value systems, cognitive frameworks, and personal experiences of annotators (sap-etal-2019-risk; truthlie; milkowski-etal-2021-personal). Recent work has introduced this concept in humans to LLMs, e.g., with socio-demographic prompting (2025demo; dayanik-etal-2022-analysis). Our approach considers the distribution of predictions by prompting LLMs a large number of times without anchoring responses to predefined demographic categories.
While subjectivity and perspectivism highlight the inherent variability in human interpretation, annotator disagreement also points to varying annotations which might not be of any benefit. Disagreement is therefore also observed in tasks considered objective, demanding a unified consensus for both model training and evaluation (disagree).
For example, plank-etal-2014-linguistically demonstrate that disagreements in part-of-speech tagging stem from linguistically debatable cases rather than annotator mistakes. Similarly, webber-joshi-2012-discourse analyze the challenges of human disagreement in discourse properties. Semantic annotation tasks are also prone to disagreement. sommerauer-etal-2020-describe evaluate such cases using multiple quality metrics. De Marneffe (didithappen) argue that fact judgments should be treated as distributions.
@@ 2.2.1 Subjectivity & Annotator Disagreement
2.2.2 Survey Response Bias
In social science studies, human responses often exhibit biases influenced by different survey designs, such as variations in wording, synonyms, and question formats used in market research (brace2008questionnaire; responseAlt). Also, the order of answer options in questionnaires can give rise to acquiescence and recency effects (responseOrd).
tjuatja-etal-2024-llms present the only study we are aware of which studies the relation between LLMs and human variations. They demonstrate that commercial LLMs exhibit sensitivity to prompt biases that elicit minimal change in human responses. However, their exploration of prompt modifications primarily relies on survey-style prompt variants and does not study human annotation variability. In contrast, our work systematically generalizes a broader and more diverse set of prompt reformulations, aiming to reflect the correlation of variability between LLMs and human annotation behavior.
@@ 3. How do you rate … given one of the labels ‘a’; ‘b’; ‘c’; ‘d’?
Table 1: Prompt modification methods in the neutral category. Modifications are bold in the text for each example. Our hypothesis for these prompt variations is that human’s show lower variation than LLMs.
@@ 3 Prompt Perturbation Methods
Sensitive changes might include the introduction of emotionally connotated wording in the prompt (referred to as Emo). For this aspect, appending the sentence “Respond quickly!” to the prompt corresponds to the variant fast. Appendix A.2 provides the full list of evaluated prompts in our experiments, developed by these methods.
@@ 4. How do you rate … given one of the labels ‘a’, ‘b’, ‘c’, ‘d’? Respond fast!
Table 2: Prompt modification methods in the sensitive category. Modifications are bold in the text for each example. We assume that humans are more affected by these variations than by those in the neutral category.
@@ 4 Experimental Settings
For both humans and LLMs, we consider four tasks from three distinct English-language datasets: (1) offensiveness rating and (2) politeness rating from the Popquorn dataset (pei-jurgens-2023-annotator), (3) irony detection from the EPIC dataset (frenda-etal-2023-epic), and (4) emotion classification from the crowd-enVent dataset (troiano-etal-2023-dimensional). These tasks were selected as all are known to involve some level of annotator subjectivity. For each dataset, we construct a base prompt, which clearly and succinctly requests an annotation for a provided instance according to the task’s label set. We then construct a set of prompt variations, as described in Section 3. A full list of all prompt variations thus obtained is detailed in Appendix A.2.
@@
We select five LLMs capable of local deployment for evaluation: LLaMA-3.1-8B and LLaMA-3.3-70B (llama3), Mixtral-8x7B (mixtral), Falcon3-7B (Falcon3), and Mistral-7B (mistral7b). [^1] https://huggingface.co/: tiiuae/Falcon3-7B-Instruct, mistralai/Mixtral-8x7B-Instruct-v0.1, mistralai/Mistral-7B-Instruct-v0.3, meta-llama/Llama-3.3-70B-Instruct, meta-llama/Llama-3.1-8B-Instruct.
@@
Model predictions are elicited in a zero-shot fashion: For each instance, the model is presented with one prompt variation, followed by the instance text. The generated text is then parsed by attempting to extract one of the requested labels, with unparsable responses discarded. Across all tasks, for each (prompt variation, instance, LLM) triple, we elicit 100 responses in this manner. We discard instances for which we obtain less than 100 total valid predictions after 500 attempts. This leaves us with a total of 9329 retained instances across all tasks, for which the statistics are shown in Table 3.
(a) Offensiveness rating.
(b) Politeness rating.
(c) Emotion classification.
(d) Irony detection.
Figure 1: Heatmaps showing distance scores between the distributions of LLM predictions with prompt variants and with the base prompt for four evaluated tasks in Table 3. Distance scores are calculated using Jensen–Shannon divergence (dagan-etal-1997-similarity). The x-axis represents different types of prompt modifications defined in Section 3, with numeric indices representing multiple instances of the same type-variant pair. The y-axis lists the five LLMs evaluated in the study. Black dashed lines divide each heatmap into three parts: prompt modifications belonging to the neutral class (left), those from the sensitive class (middle), and the average distance scores for neutral and sensitive class modifications respectively (right). All prompts evaluated are provided in Appendix A.2.
@@
Figure 2: Distance scores between the distributions of responses with prompt variants and the base prompt. The evaluated tasks are offensiveness rating (a) and emotion classification (b). The x-axis refers to the prompts with their modification types, and the y-axis refers to five LLMs and human samples. Distance scores are measured by Jensen–Shannon divergence (dagan-etal-1997-similarity). Evaluated prompts can be found in Table LABEL:tab:prompt-off (a) and Table LABEL:tab:prompt-emo (b).
Human Study.
Due to the high cost of human annotation, we collect a subset of the judgments from our LLM study for human evaluation, focusing on two tasks: offensiveness rating and emotion classification. For each task, we sample approximately 10% of instances from the original Popquorn and crowd-enVent datasets, selecting an equal number of instances for each gold-standard label. We use ten prompts per task for the human study, selected so as to maximize the diversity of prompts tested. In addition to the base prompt, we select six variants which caused the most, medium and least extreme changes in LLM predictions from both the sensitive and neutral categories, as measured by Jensen–Shannon divergence (see Section 5). The other three variants are selected from the remaining types which present notable changes in instructing humans. Each prompt-instance pair is annotated by ten independent annotators recruited via the Prolific platform. [^2] https://www.prolific.com/ We employ only annotators based in the United States, with U.S. listed as their country of residence, country of birth, and nationality. All annotators are at least 18 years of age and demonstrated a high level of English proficiency, with English specified as their primary, first, and fluent language. We restrict the annotator approval rating to greater than 99%. Each annotator is presented with 22 instances including two attention checks, each preceded by a prompt variation acting as instructions. To mitigate potential biases arising from prior exposure to the same task, each annotator is assigned only one prompt variant per task. Each annotator is paid £9 per hour as suggested by Prolific. The median completion time for each survey is approximately 7–8 minutes, except for the Exp prompt variant, which requires about 16 minutes. The whole experiment costs £2,400, including 33.3% Prolific service fee.